# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 2.2. Stars, mixed-tree-decompositions and torsos
The following definition is supported by Figure 2.2. Let $G$ be a graph. A pair $\mathcal{T}=(T,(V_{t})_{t\in T})$ of a tree $T$ and a family of vertex sets $V_{t}\subseteq V(G)$ indexed by the nodes $t\in T$ is a mixed-tree-decomposition of $G$ if it satisfies the following two conditions:
@@
Figure 2.2. A mixed-tree-decomposition. The decomposition-tree is blue. Bags are grey. Edges and vertices in adhesion-sets are red.
A star of mixed-separations+, or star for short, is a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations+ $(A_{i},B_{i})$ such that $(A_{i},B_{i})\leqslant(B_{j},A_{j})$ for all distinct $i,j\in I$. For every node $t$ of $T$, the set $\sigma_{t}:=\{\,\alpha_{\mathcal{T}}(s,t):st\in E(T)\,\}$ is a star. Now let $S$ be a set of mixed-separations+ with $\sigma\subseteq S$. We call $\sigma$ a splitting star of $S$ if $\sigma$ is a star and for every $(C,D)\in S$ there is $i\in I$ with either $(C,D)\leqslant(A_{i},B_{i})$ or $(D,C)\leqslant(A_{i},B_{i})$. If $\mathcal{T}$ is proper, then the splitting stars of the image of $\alpha_{\mathcal{T}}$ are precisely the stars $\sigma_{t}$. Conversely, if $S$ is a nested symmetric set of mixed-separations of a finite graph $G$, then there is a unique (up to isomorphisms for decomposition-trees) mixed-tree-decomposition $\mathcal{T}(S)$ of $G$ such that $\alpha_{\mathcal{T}(S)}$ is a bijection $\vec{E}(T)\to S$. The mixed-tree-decomposition $\mathcal{T}(S)=(T,\mathcal{V})$ has an explicit description (see [Tridecomp, §3.1]):
@@
When $G$ is infinite and $S$ is the set of totally-nested tetra-separations of $G$, then $\mathcal{T}(S)$ is a mixed-tree-decomposition of $G$ by standard arguments as in [Tridecomp, Lemma 3.3.10].
A mixed-separation $(C,D)$ interlaces a star $\sigma$ of mixed-separations if for every $(A,B)\in\sigma$ we have either $(A,B)<(C,D)$ or $(A,B)<(D,C)$. The notion of interlacing yields a characterisation of splitting stars: a star $\sigma$ included in a nested set $S$ of mixed-separations of a graph $G$ is a splitting star of $S$ if and only if no element of $S$ interlaces $\sigma$ [Tridecomp, Lemma 2.2.3].
Let $\mathcal{T}=(T,\mathcal{V})$ be a mixed-tree-decomposition of a graph $G$. For a mixed-separation+ $(A,B)$, we let the vertex-set $)𝜆𝐴𝐵{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{\lambda(A,B)}}$ consist of all vertices in $S(A,B)$ and the endvertex in $A\smallsetminus B$ for each edge in $S(A,B)$. The expanded-torso of $t$, or of $V_{t}$, is obtained from the subgraph of $G$ induced by
@@
by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique. From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure 2.3.
@@
Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure 2.2.
@@ 2.3. Graph-decompositions
Let $X$ be a graph. We call $X$ a generalised wheel with centre $v$ if $X$ is 3-connected with $v\in V(X)$ such that $X-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size two such that every torso is isomorphic to some $X_{i}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
3. Proofs of Theorem 1 and 2 and strengthenings thereof
Instead of proving Theorem 1 and 2 directly, we will first prove stronger but slightly more technical versions of them, Theorem 3.1 and Corollary 3.2 below, and then immediately obtain Theorem 1 and 2 from their strengthenings. The proof of Theorem 3.1 uses a number of lemmas that have not yet been introduced. Stating and proving these lemmas will be the main objectives of later sections.
Theorem 3.1 (Technical strengthening of Theorem 1).
@@ 3. Proofs of Theorem 1 and 2 and strengthenings thereof
If no tetra-separation of $G$ interlaces $\sigma$, then $\tau$ is quasi-5-connected by 10.5. So we may assume that some tetra-separation $(A,B)$ of $G$ interlaces $\sigma$. Then $(A,B)$ is crossed by some tetra-separation $(C,D)$, with $Z$ denoting the vertex-centre. By the (Crossing Lemma). (4.1), we have $|Z|\in\{0,2,4\}$.
(ii). Assume $|Z|=0$. Then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices by 9.50.
(iii). Assume $|Z|=2$. Then $\tau$ is a generalised double-wheel with centre $Z$ by 8.37.
(iv). Assume $|Z|=4$. If $\sigma$ is non-empty, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as left side and $m\geqslant 0$ by 7.7. Otherwise the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$, also by 7.7. ∎
Proof of Theorem 1.
Immediate from Theorem 3.1. ∎
Corollary 3.2 (Technical strengthening of 2).
@@
(i). Assume that $\sigma$ is not interlaced by any strict tri-separation. Then $\hat{\sigma}$ is not interlaced by any tetra-separation, so $\hat{\tau}$ is quasi-5-connected by Theorem 3.1 (i). If $\tau$ is not quasi-4-connected, this is witnessed by some separation $(X,Y)$ of $\tau$, and then $(X\cup\{\alpha\},Y\cup\{\alpha\})$ witnesses that $\hat{\tau}$ is not quasi-5-connected. Hence $\tau$ is quasi-4-connected.
@@
(ii). Assume that $Z$ consists of exactly one vertex $v$. Then $\hat{\tau}$ is a generalised double-wheel with centre $\{v,\alpha\}$ by Theorem 3.1 (iii). Hence $\hat{\tau}$ is a generalised wheel with centre $v$.
(iii). The vertex-centre $Z$ cannot have sizes 0 or 2, since then $\hat{Z}$ would have size 1 or 3, which is excluded in Theorem 3.1. Thus $|Z|=3$, so $|\hat{Z}|=4$. Since $\alpha\in\hat{Z}$, both outcomes of Theorem 3.1 (iv) translate to the two outcomes of (iii). ∎
Proof of 2.
Immediate from Corollary 3.2. ∎
@@
The outcome in Theorem 3.1 is not in general unique. For example, consider the graph $G$ depicted in Figure 3.1 on the left. Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$. On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem 3.1 (ii). On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2. From their perspective, $G$ is a generalised wheel, see Theorem 3.1 (iii).
For another example, consider the graph $G$ depicted in Figure 3.2 on the left. Again, $\sigma=\emptyset$ is a splitting star of $N(G)$. On the left of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem 3.1 (ii). On the right of Figure 3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4. From their perspective, $G$ is a $K_{4,4}$, see Theorem 3.1 (iv).
@@ 4. How tetra-separations cross
We require the following terminology for corner diagrams from [Tridecomp, §1.3], which is supported by Figure 4.1. Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. All of the following definitions depend on $(A,B)$ and $(C,D)$ being clear from context. The corner for $\{A,C\}$, or the corner $AC$ for short, is the vertex $(A\smallsetminus B)\cap(C\smallsetminus D)$. The corners for the other three combinations of one side from $(A,B)$ and one side from $(C,D)$ are defined analogously. Two corners are adjacent when their pair share a side, and opposite otherwise. An edge of $G$ is diagonal if its endvertices lie in opposite corners. The centre consists of all vertices in the vertex-centre and all diagonal edges.
@@
The following definition is supported by Figure 4.2. The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
[Tridecomp, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
@@
Otherwise, two adjacent corner-separators have size at most 3, say the corner-separator for $AC$ and the corner-separator for $BC$. By 4-connectivity, the corners for $AC$ and $BC$ are empty. Since $C\smallsetminus D\neq\emptyset$, the link $L$ for $C$ is nonempty. The link $L$ only contains vertices since its adjacent corners are empty, so $L=(A\cap B)\smallsetminus D$. If $|L|\geqslant 2$ then the vertices in $L$ have at most one neighbour in $A\smallsetminus B$ (and $B\smallsetminus A$), contradicting the degree-condition for $(A,B)$. Otherwise, if $|L|=1$ then $(C,D)$ is not a tetra-separation by Lemma 4.2, a contradiction. ∎
The following lemma may be regarded as an analogue of [Tridecomp, Lemma 1.3.5].
@@
    two adjacent links have size $i$, the other two links have size $j$, and the centre has size $k-i-j$, where $1\leqslant i\leqslant\lfloor(k-1)/2\rfloor$ and $i<j\leqslant k-i$.
@@
By Lemma 4.3, two opposite corner-separators have size exactly $k$, say the corner-separators for $AC$ and $BD$. Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively. Let $x$ denote the size of the centre. By Lemma 4.3, every diagonal edge has its endvertices in the corners for $AC$ and $BD$; and there are no jumping edges. We get $a+c+x=k$ and $b+d+x=k$ and $a+b+x=k$ and $c+d+x=k$. It follows that $b=c$ and $a=d$. Without loss of generality $a=d\leqslant b=c$.
@@
Otherwise $1\leqslant a=d\leqslant\lfloor k/2\rfloor$. If $a=b=c=d=:\ell$ then we get $x=k-a-c=k-2\ell$ and so get outcome (i). Otherwise $i:=a=d<b=c=:j$. Note that $i\leqslant\lfloor(k-1)/2\rfloor$. With $x=k-a-c=k-i-j$, we get outcome (ii). ∎
@@
Suppose that not all links have size $\ell$, for some $\ell\in\{0,1,2\}$. By Lemma 4.4 and Lemma 4.5, we get that two adjacent links have size one, the other two links have size at least two and the centre has size at most one. Without loss of generality, the corner $K$ of $AC$ is the corner whose adjacent links have size one. The corner-separator of $K$ has size at most 3, so $K$ is empty since $G$ is 4-connected. Then one of the links of size one contains a vertex $u$, say $u\in(A\cap B)\smallsetminus D$. But then $u$ has at most one neighbour in $A\smallsetminus B$, contradicting the degree-condition for $(A,B)$. Hence all links have the same size $\ell$ for some $\ell\in\{0,1,2\}$, and then the centre has size $4-2\ell$ by Lemma 4.4 and Lemma 4.5. ∎
@@
(i) By Lemma 4.6, all links have the same size $\ell$ and the centre has size $4-2\ell$. By Lemma 4.3, there do not exist jumping edges nor diagonal edges.
(ii) Without loss of generality, $X=A$ and $Y=C$. Let $e=u_{1}u_{2}$ be a dangling edge such that $u_{2}$ lies in the $C$-link and $e$ lies in the $A$-link. We first show that the corner $K$ for $AC$ is empty. Assume for contradiction that $K\neq\emptyset$. By Lemma 4.6 and 4-connectivity, $|L(A,C)|=4$. Then $L(A,C)-e$ is a mixed-3-separator separating the vertices in the corner for $AC$ from $u_{1}$. This contradicts 4-connectivity by Lemma 6.1.
By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{2}$ in the $C$-link has a neighbour $u_{3}$ in the $A$-link. By the degree-condition of $(C,D)$ and since $K$ is empty, the vertex $u_{3}$ has a neighbour $u_{4}$ in $(C\smallsetminus D)\cap B$. It remains to show that $u_{4}$ is in the corner for $BC$. Otherwise, $u_{4}$ is in the $C$-link. By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{4}$ has a neighbour in $A\cap D$ besides $u_{3}$, and therefore the $A$-link contains a third vertex or edge besides $e$ and $u_{3}$. This contradicts Lemma 4.6. ∎
@@
We combine Lemma 4.6 and Lemma 4.8. ∎
5. Shifts: a tool for finding tetra-separations
This section provides a ‘shifting’ method for obtaining tetra-separations from genuine 4-separations, and from mixed-4-separations more generally. There are ‘small’ mixed-4-separations that fail to yield tetra-separations via the shifting method. We characterise these in 5.10.
5.1. Left-shifts and right-shifts
Definition 5.1 (Left-shift, right-shift).
Let $(A,B)$ be a mixed-separation+ in a graph. Let $L$ be the set of vertices in $A\cap B$ with at most one neighbour in $A\smallsetminus B$, and let $R$ be the set of vertices in $A\cap B$ with at most one neighbour in $B\smallsetminus A$. The left-shift of $(A,B)$ is $(A\smallsetminus L,B)$. The right-shift of $(A,B)$ is $(A,B\smallsetminus R)$.
In the following, we prove lemmas for the left-shift. All results have analogues for right-shifts by symmetry.

### v2

@@ 2.2. Stars, mixed-tree-decompositions and torsos
The following definition is supported by Figure˜2.2. Let $G$ be a graph. A pair $\mathcal{T}=(T,(V_{t})_{t\in T})$ of a tree $T$ and a family of vertex sets $V_{t}\subseteq V(G)$ indexed by the nodes $t\in T$ is a mixed-tree-decomposition of $G$ if it satisfies the following two conditions:
@@
Figure $2.$ A mixed-tree-decomposition. The decomposition-tree is blue. Bags are grey. Edges and vertices in adhesion-sets are red.
A star of mixed-separations+, or star for short, is a set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations+ $(A_{i},B_{i})$ such that $(A_{i},B_{i})\leqslant(B_{j},A_{j})$ for all distinct $i,j\in I$. For every node $t$ of $T$, the set $\sigma_{t}:=\{\,\alpha_{\mathcal{T}}(s,t):st\in E(T)\,\}$ is a star. Now let $S$ be a set of mixed-separations+ with $\sigma\subseteq S$. We call $\sigma$ a splitting star of $S$ if $\sigma$ is a star and for every $(C,D)\in S$ there is $i\in I$ with either $(C,D)\leqslant(A_{i},B_{i})$ or $(D,C)\leqslant(A_{i},B_{i})$. If $\mathcal{T}$ is proper, then the splitting stars of the image of $\alpha_{\mathcal{T}}$ are precisely the stars $\sigma_{t}$. Conversely, if $S$ is a nested symmetric set of mixed-separations of a finite graph $G$, then there is a unique (up to isomorphisms for decomposition-trees) mixed-tree-decomposition $\mathcal{T}(S)$ of $G$ such that $\alpha_{\mathcal{T}(S)}$ is a bijection $\vec{E}(T)\to S$. The mixed-tree-decomposition $\mathcal{T}(S)=(T,\mathcal{V})$ has an explicit description (see [CK23, §3.1]):
@@
When $G$ is infinite and $S$ is the set of totally-nested tetra-separations of $G$, then $\mathcal{T}(S)$ is a mixed-tree-decomposition of $G$ by standard arguments as in [CK23, Lemma 3.3.10].
A mixed-separation $(C,D)$ interlaces a star $\sigma$ of mixed-separations if for every $(A,B)\in\sigma$ we have either $(A,B)<(C,D)$ or $(A,B)<(D,C)$. The notion of interlacing yields a characterisation of splitting stars: a star $\sigma$ included in a nested set $S$ of mixed-separations of a graph $G$ is a splitting star of $S$ if and only if no element of $S$ interlaces $\sigma$ [CK23, Lemma 2.2.3].
[…]
@@
by making each of the vertex-sets $\lambda(\alpha_{\mathcal{T}}(s,t))$ into a clique. From this expanded-torso we obtain the compressed-torso, or torso for short, by contracting all edges contained in adhesion-sets of the edges $st\in E(T)$; see Figure˜2.3.
@@
Figure 2.3. The torsos of the mixed-tree-decomposition depicted in Figure˜$2.$
@@ 2.3. Graph-decompositions
Let $X$ be a graph. We call $X$ a generalised wheel with centre $v$ if $X$ is 3-connected with $v\in V(X)$ such that XvX-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a generalised double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition with all adhesion-sets of size 1 and into $K_{2}$’s and triangles. We call $X$ a cycle of $X_{1}$-torsos, …, $X_{k}$-torsos where the $X_{i}$ are graphs, if $X$ is 4-connected and has a cycle-decomposition with all adhesion-sets of size two such that every torso is isomorphic to some $X_{i}$. A thickened $K_{4,m}$ is obtained from a $K_{4,m}$ by turning the left side into a $K_{4}$. A sprinkled $K_{4,m}$ is obtained from a $K_{4,m}$ by adding any number of edges to the left side.
3. Proofs of Theorem˜1 and ˜2 and strengthenings thereof
Instead of proving Theorem˜1 and ˜2 directly, we will first prove stronger but slightly more technical versions of them, Theorem˜3.1 and Corollary˜3.2 below, and then immediately obtain Theorem˜1 and ˜2 from their strengthenings. The proof of Theorem˜3.1 uses a number of lemmas that have not yet been introduced. Stating and proving these lemmas will be the main objectives of later sections.
Theorem 3.1 (Technical strengthening of Theorem˜1).
@@ 3. Proofs of Theorem˜1 and ˜2 and strengthenings thereof
If no tetra-separation of $G$ interlaces $\sigma$, then $\tau$ is quasi-5-connected by ˜10.5. So we may assume that some tetra-separation $(A,B)$ of $G$ interlaces $\sigma$. Then $(A,B)$ is crossed by some tetra-separation $(C,D)$, with $Z$ denoting the vertex-centre. By the (Crossing Lemma). (4.1), we have $|Z|\in\{0,2,4\}$.
˜(ii). Assume $|Z|=0$. Then $\tau$ is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices by ˜9.50.
˜(iii). Assume $|Z|=2$. Then $\tau$ is a generalised double-wheel with centre $Z$ by ˜8.37.
˜(iv). Assume $|Z|=4$. If $\sigma$ is non-empty, then $\tau$ is a thickened $K_{4,m}$ with $Z$ as left side and $m\geqslant 0$ by ˜$7.$ Otherwise the entire graph $G$ is a sprinkled $K_{4,m}$ with $Z$ as left side and $m\geqslant 4$, also by ˜$7.$ ∎
Proof of Theorem˜1.
Immediate from Theorem˜3.1. ∎
Corollary 3.2 (Technical strengthening of ˜2).
@@
˜(i). Assume that $\sigma$ is not interlaced by any strict tri-separation. Then $\hat{\sigma}$ is not interlaced by any tetra-separation, so $\hat{\tau}$ is quasi-5-connected by Theorem˜3.1 ˜(i). If $\tau$ is not quasi-4-connected, this is witnessed by some separation $(X,Y)$ of $\tau$, and then $(X\cup\{\alpha\},Y\cup\{\alpha\})$ witnesses that $\hat{\tau}$ is not quasi-5-connected. Hence $\tau$ is quasi-4-connected.
@@
˜(ii). Assume that $Z$ consists of exactly one vertex $v$. Then $\hat{\tau}$ is a generalised double-wheel with centre $\{v,\alpha\}$ by Theorem˜3.1 ˜(iii). Hence $\hat{\tau}$ is a generalised wheel with centre $v$.
˜(iii). The vertex-centre $Z$ cannot have sizes 0 or 2, since then $\hat{Z}$ would have size 1 or 3, which is excluded in Theorem˜3.1. Thus $|Z|=3$, so $|\hat{Z}|=4$. Since $\alpha\in\hat{Z}$, both outcomes of Theorem˜3.1 ˜(iv) translate to the two outcomes of ˜(iii). ∎
Proof of ˜2.
Immediate from Corollary˜3.2. ∎
@@
The outcome in Theorem˜3.1 is not in general unique. For example, consider the graph $G$ depicted in Figure˜3.1 on the left. Every tetra-separation of $G$ is crossed, hence $\sigma=\emptyset$ is a splitting star of the set $N(G)$ of all totally-nested tetra-separations of $G$. On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos and $4$-wheel torsos, see Theorem˜3.1 ˜(ii). On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 2. From their perspective, $G$ is a generalised wheel, see Theorem˜3.1 ˜(iii).
For another example, consider the graph $G$ depicted in Figure˜3.2 on the left. Again, $\sigma=\emptyset$ is a splitting star of $N(G)$. On the left of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with empty vertex-centre. From their perspective, $G$ is a cycle of $K_{4}$-torsos, see Theorem˜3.1 ˜(ii). On the right of Figure˜3.1, we see two tetra-separations that interlace $\sigma$ and cross with a vertex-centre of size 4. From their perspective, $G$ is a $K_{4,4}$, see Theorem˜3.1 ˜(iv).
@@ 4. How tetra-separations cross
We require the following terminology for corner diagrams from [CK23, §1.3], which is supported by Figure˜4.1. Let $(A,B)$ and $(C,D)$ be mixed-separations+ of a graph $G$. All of the following definitions depend on $(A,B)$ and $(C,D)$ being clear from context. The corner for $\{A,C\}$, or the corner $AC$ for short, is the vertex $(A\smallsetminus B)\cap(C\smallsetminus D)$. The corners for the other three combinations of one side from $(A,B)$ and one side from $(C,D)$ are defined analogously. Two corners are adjacent when their pair share a side, and opposite otherwise. An edge of $G$ is diagonal if its endvertices lie in opposite corners. The centre consists of all vertices in the vertex-centre and all diagonal edges.
@@
The following definition is supported by Figure˜4.2. The corner for $AC$ is potter if it is empty and its adjacent links have size two and there exists a path $u_{1}u_{2}u_{3}u_{4}$ in $G$ such that $u_{1}u_{2}$ dangles from the vertex $u_{2}$ in the $C$-link through the $A$-link, and $u_{3}u_{4}$ dangles from the vertex $u_{3}$ in the $A$-link through the $C$-link.
@@
[CK23, Lemma 1.3.4] For two mixed-separations $(A,B)$ and $(C,D)$ of $G$, we have
@@
Otherwise, two adjacent corner-separators have size at most 3, say the corner-separator for $AC$ and the corner-separator for $BC$. By 4-connectivity, the corners for $AC$ and $BC$ are empty. Since $C\smallsetminus D\neq\emptyset$, the link $L$ for $C$ is nonempty. The link $L$ only contains vertices since its adjacent corners are empty, so $L=(A\cap B)\smallsetminus D$. If $|L|\geqslant 2$ then the vertices in $L$ have at most one neighbour in $A\smallsetminus B$ (and $B\smallsetminus A$), contradicting the degree-condition for $(A,B)$. Otherwise, if $|L|=1$ then $(C,D)$ is not a tetra-separation by Lemma˜4.2, a contradiction. ∎
The following lemma may be regarded as an analogue of [CK23, Lemma 1.3.5].
@@
    two adjacent links have size $i$, the other two links have size $j$, and the centre has size kijk-i-j, where $1\leqslant i\leqslant\lfloor(k-1)/2\rfloor$ and $i<j\leqslant k-i$.
@@
By Lemma˜4.3, two opposite corner-separators have size exactly $k$, say the corner-separators for $AC$ and $BD$. Let $a,b,c,d$ denote the sizes of the links for $A,B,C,D$, respectively. Let $x$ denote the size of the centre. By Lemma˜4.3, every diagonal edge has its endvertices in the corners for $AC$ and $BD$; and there are no jumping edges. We get $a+c+x=k$ and $b+d+x=k$ and $a+b+x=k$ and $c+d+x=k$. It follows that $b=c$ and $a=d$. Without loss of generality $a=d\leqslant b=c$.
@@
Otherwise $1\leqslant a=d\leqslant\lfloor k/2\rfloor$. If $a=b=c=d=:\ell$ then we get $x=k-a-c=k-2\ell$ and so get outcome (i). Otherwise $i:=a=d<b=c=:j$. Note that $i\leqslant\lfloor(k-1)/2\rfloor$. With x=kac=kijx=k-a-c=k-i-j, we get outcome (ii). ∎
@@
Suppose that not all links have size $\ell$, for some $\ell\in\{0,1,2\}$. By Lemma˜4.4 and Lemma˜4.5, we get that two adjacent links have size one, the other two links have size at least two and the centre has size at most one. Without loss of generality, the corner $K$ of $AC$ is the corner whose adjacent links have size one. The corner-separator of $K$ has size at most 3, so $K$ is empty since $G$ is 4-connected. Then one of the links of size one contains a vertex $u$, say $u\in(A\cap B)\smallsetminus D$. But then $u$ has at most one neighbour in $A\smallsetminus B$, contradicting the degree-condition for $(A,B)$. Hence all links have the same size $\ell$ for some $\ell\in\{0,1,2\}$, and then the centre has size $4-2\ell$ by Lemma˜4.4 and Lemma˜4.5. ∎
@@
(i) By Lemma˜4.6, all links have the same size $\ell$ and the centre has size $4-2\ell$. By Lemma˜4.3, there do not exist jumping edges nor diagonal edges.
(ii) Without loss of generality, $X=A$ and $Y=C$. Let $e=u_{1}u_{2}$ be a dangling edge such that $u_{2}$ lies in the $C$-link and $e$ lies in the $A$-link. We first show that the corner $K$ for $AC$ is empty. Assume for contradiction that $K\neq\emptyset$. By Lemma˜4.6 and 4-connectivity, $|L(A,C)|=4$. Then L(A,C)eL(A,C)-e is a mixed-3-separator separating the vertices in the corner for $AC$ from $u_{1}$. This contradicts 4-connectivity by Lemma˜6.1.
By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{2}$ in the $C$-link has a neighbour $u_{3}$ in the $A$-link. By the degree-condition of $(C,D)$ and since $K$ is empty, the vertex $u_{3}$ has a neighbour $u_{4}$ in $(C\smallsetminus D)\cap B$. It remains to show that $u_{4}$ is in the corner for $BC$. Otherwise, $u_{4}$ is in the $C$-link. By the degree-condition of $(A,B)$ and since $K$ is empty, the vertex $u_{4}$ has a neighbour in $A\cap D$ besides $u_{3}$, and therefore the $A$-link contains a third vertex or edge besides $e$ and $u_{3}$. This contradicts Lemma˜4.6. ∎
@@
We combine Lemma˜4.6 and Lemma˜4.8. ∎
5. Reductions: a tool for finding tetra-separations
This section provides a ‘reduction’ method for obtaining tetra-separations from genuine 4-separations, and from mixed-4-separations more generally. There are ‘small’ mixed-4-separations that fail to yield tetra-separations via the reduction method. We characterise these in ˜5.10.
5.1. Left-reductions and right-reductions
Definition 5.1 (Left-reduction, right-reduction).
Let $(A,B)$ be a mixed-separation+ in a graph. Let $L$ be the set of vertices in $A\cap B$ with at most one neighbour in $A\smallsetminus B$, and let $R$ be the set of vertices in $A\cap B$ with at most one neighbour in $B\smallsetminus A$. The left-reduction of $(A,B)$ is $(A\smallsetminus L,B)$. The right-reduction of $(A,B)$ is $(A,B\smallsetminus R)$.
In the following, we prove lemmas for the left-reduction. All results have analogues for right-reductions by symmetry.